# The Gorini-Kossakowski-Sudarshan-Lindblad problem and the geometry of CP maps

### v3

@@ 1 Introduction
§5 turns the development toward the GKSL problem. We are interested CP families of evolution operators $\mathcal{E}(t,s)$ for $t\geq s$ satisfying the differential equation
@@
and the Markovianity condition $u\geq t\geq s\Rightarrow\mathcal{E}(u,s)=\mathcal{E}(u,t)\circ\mathcal{E}(t,s)$. The question is one of necessary and sufficient conditions on the generator $\mathscr{L}(t)$. It is not hard to see, first of all, that the generator must at least lie in the tangent cone to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$, denoted $\mathsf{cp}^{+}(\mathcal{H})$. The relevant ambient vector space needed to make sense of tangency here is $\mathsf{HP}(\mathcal{H})$, consisting of hermiticity-preserving superoperators, and generated by the closed convex cone $\mathsf{CP}(\mathcal{H})$. Then it is shown that the map $L\colon\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\rightarrow\mathsf{HP}(\mathcal{H})$ ($\mathrm{SA}(\mathcal{H})$ is the space of hermitian operators) given by
$L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;]$ (1.5)
yields acceptable generators so long as $\Psi$ is CP. The $L(\Psi,G,H)$ for such $\Psi$ comprise the “$L$-cone”. An operator varying continuously in the $L$-cone generates a Markovian CP evolution. The proof of this given in §5.4 works in both finite and infinite dimensions; the latter case merely requires that we specify continuity in operator norm.
The treatment of the “necessity” direction given in §5 uses the Jamiołkowski transform and is therefore limited to finite dimensions. §5.7 shows that there exist linear right-inverses $\Delta\colon\mathsf{HP}(\mathcal{H})\rightarrow\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ to $L$ (i.e., $L\circ\Delta=\operatorname{\mathbbm{1}}_{\mathsf{HP}(\mathcal{H})}$) which map the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ into $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$. The choice of such a Lindblad parametrizer amounts to a choice of complementary subspace to the tangent space $\mathsf{cp}(\mathcal{H})$ to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$. Thus, the $L$-cone is all of $\mathsf{cp}^{+}(\mathcal{H})$, and uniform ways of extracting a $(\Psi,G,H)$ parameterization exist. This not only clarifies the situation relative to a muddy can-be-written claim, but is of importance when dealing with time-varying generators or perturbation problems.
The second part of the paper, starting in §7, is devoted to handling the infinite-dimensional (separable) case, bootstrapping from the finite-dimensional theory by way of a sequence of finite-dimensional approximations. With that approach, instead of the operator-algebraic methods used by Kraus[11, 12] to show existence of Kraus decompositions and by Lindblad[16] for the GKSL problem, we manage with more basic functional analytic tools.
The price to be paid is that we need some “home-made” tools. After a brief orientation toward the separable setting in §7, these are developed in §§8 and 9. The basic idea is an increasing sequences of finite-dimensional approximations, dubbed “filtration”. §10 shows how these tools are useful in studying the geometry of CP maps. Just as in finite dimensions, extreme vectors in the convex CP cone are those of form $\rho\mapsto A\rho A^{\dagger}$, a map we denote ${\bm{\theta}}(A)$. To obtain Kraus decomposition in this approach, we show that we can subtract a map ${\bm{\theta}}(A)$ from any nonzero CP map with a remainder which is still CP. Then we simply repeat this. With some good choices of $A$, convergence can be shown. Even given the finite-dimensional result, the existence of Kraus decompositions in the separable setting is prima facie surprising. From a convex analysis perspective, or considering the phenomenon of continuous spectrum, some sort of integral rather than a simple sum might well be expected. Finally, §11 tackles the GKSL problem in the separable setting, recovering analogs of all the finite-dimensional results. A handy table of notations and terminology can be found in Appendix A.
@@ 2.3 Jamiołkowski transformation
2.3.1 A fundamental isomorphism
@@
2.3.2 Jamiołkowski transformation
@@ 3.2.1 The trouble with $\mathsf{Mon}$
3.2.2 Making the world safe for tensor products
@@ 7.4 Meanings of old notations in the separable setting
$\mathsf{Lind\,{\mathcal{H}}}=\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H}),$ (7.9)
@@
8 Filtrations
@@ 8.3 Generic notation
9 The $d$-metric
@@
The $d$-topology is very weak. For instance, for $T\in\mathfrak{B}(\mathcal{H},\mathcal{K})$, $\mathscr{P}_{\alpha}T\mathscr{P}_{\alpha}$ amounts to a finite number of coordinates (linear functionals) of $Tv$ for a finite number of $v$ (basis of $\mathcal{H}_{\alpha}$). Thus, it is weaker than WOT. To put it another way, there are much fewer $d$-closed (or open) sets that WOT-closed, in some vague sense. Thus, it is gratifying and useful to have the following criterion.
@@ 10 CP maps in separable setting
The equivalence of item 1 above has its own interest, but is also crucial to item 2. The standard proof that Kraus decompositions exists rests on Stinespring’s theorem and a knowledge of the representations of the $C^{*}$-algebra $\mathfrak{B}(\mathcal{H})$. An explanation of how to use these results to show existence of Kraus decompositions can be found in Theorem 3.3 of [11] or Chapter 3 of [12]. Expositions of Stinespring’s theorem itself can be found in 34.7 of Ref. [5], Thm. IX.4.2 of Ref. [6], Thm. IV.3.6 of Ref. [26].
In comparison to the traditional method, which feels like a deus ex machina, the approach taken here is very pedestrian. We show that for any nonzero CP map $\Lambda$, we can subtract a $\bm{\theta}$}-map and have a CP map left over. Then we show that if one just keeps doing that (with some good choices), everything will converge as desired.
@@
10.1 Partial orders on hermiticity-preserving maps
@@
10.2 Topological closedness
@@
All this means is that if $\Lambda_{n}\precsim\Gamma_{n}$, $\Lambda_{n}\xrightarrow{d}\Lambda$, and $\Gamma_{n}\xrightarrow{d}\Gamma$, then $\Lambda\precsim\Gamma$. It follows immediately from Prop. 10.5 because what those relations mean is $\Gamma_{n}-\Lambda_{n}\in\mathsf{CP}(\mathcal{H},\mathcal{K})$.
@@ 10.4 Extreme CP map = $\bm{\theta}$-map
Let $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ with $\mathcal{H}$, $\mathcal{K}$ finite-dimensional, and let $\Lambda{\bm{\pi}}(h)$ have orthogonal decomposition $\Lambda{\bm{\pi}}(h)={\bm{\pi}}(k)+\cdots$. Then there exists a choice of $A_{i}$ in the Kraus decomposition $\Lambda=\sum_{i}{\bm{\theta}}(A_{i})$ such that
@@ 10.5 Order closedness of $\mathsf{CP}(\mathcal{H},\mathcal{K})$
Let $D$ be a $\leq$-directed set in $\mathsf{HP}(\mathcal{H},\mathcal{K})$, bounded above. Then $\sup D$ exists and $D$ converges to it in SOT. An analogous statement holds for a set directed the other way ($\geq$).
@@
If $D$ is a bounded-above $\leq$-directed set contained in $\mathsf{CP}(\mathcal{H},\mathcal{K})$, then so is its supremum. An analogous statement holds for a set directed the other way ($\geq$).
@@ 11.2 Projector juggling
$L\colon\mathsf{Lind\,{\mathcal{H}}}\rightarrow\mathsf{HP}(\mathcal{H})$ is defined by $L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;]$, just as in finite dimensions. Since the formula is just the same $(L\circ\Delta^{\scriptstyle{min}}_{\alpha})\Gamma=\Gamma$ if $\Gamma\in\mathsf{HP}(\mathcal{H}_{\alpha})$. So, if we replace $\Gamma$ in the preceding equation by $\mathscr{P}_{\alpha}\Lambda$, we get
@@
11.3 Lindblad parametrizations and parametrizers
@@ 11.3.1 The sequence $(\Delta^{\scriptstyle{min}}_{\alpha})$ is an approximate right-inverse of $L$
$L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)\xrightarrow{SOT}\Lambda$
@@
We will need the following calculation:
@@
Now, since $L$ and $\Delta^{\scriptstyle{min}}_{\alpha}$ are bounded uniformly in $\alpha$, according to Lemma 11.1, it suffices to show that $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)\cdot A\rightarrow\Lambda A$ for $A$ in the set $\cup_{m}\mathcal{B}^{1}(\mathcal{H}_{m})$ dense in $\mathcal{B}^{1}(\mathcal{H})$. Thus, take AA in $\mathcal{B}^{1}(\mathcal{H}_{\beta})$. As soon as $\beta\leq\alpha$, $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)A=(\mathscr{P}_{\alpha}\Lambda)A$, by (11.10). But, $\mathscr{P}_{\alpha}\Lambda\xrightarrow{SOT}\Lambda$, and the conclusion follows immediately. ∎
@@ 11.3.2 Bringing in the $d$-metrics
Eventually, we must work with the $d$-metrics because that is the only topology we have with respect to which any bounded sequence has cluster points.

### v4

@@ 1 Introduction
§5 turns the development toward the GKSL problem. We are interested in CP families of evolution operators $\mathcal{E}(t,s)$ for $t\geq s$ satisfying the differential equation
@@
and the Markovianity condition $u\geq t\geq s\Rightarrow\mathcal{E}(u,s)=\mathcal{E}(u,t)\circ\mathcal{E}(t,s)$. The problem is to find necessary and sufficient conditions on the generator $\mathscr{L}(t)$. It is easy to see that the generator must at least lie in the tangent cone to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$, denoted $\mathsf{cp}^{+}(\mathcal{H})$. We will show that it is sufficient if $\mathscr{L}(t)$ varies continuously in this tangent cone. Moreover, $\mathsf{cp}^{+}(\mathcal{H})$ coincides with the image (“$L$-cone”) of $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ under the linear map
$L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;],$ (1.5)
and there are linear maps $\Delta$ (“Lindblad parametrizers”) which are right inverse to $L$; and such that the first component of $\Delta\Lambda$ is CP if (and only if) $\Lambda$ is in the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ and zero if $\Lambda$ is in the tangent space $\mathsf{cp}(\mathcal{H})$. This latter space consists of the superoperators generating reversible CP evolutions.
Although both Kraus decompositions and Lindblad parametrizations both involve can-be-written type claims, Lindblad parametrizers show that the latter is in a sense much simpler as there are uniform ways of extracting a suitable $(\Psi,G,H)$ parametrization. This is of importance when dealing with time-varying generators or perturbation problems.
The second part of the paper, starting in §7, is devoted to handling the infinite-dimensional (separable) case, bootstrapping from the finite-dimensional theory by way of a sequence of finite-dimensional approximations. After a brief orientation toward the separable setting in §7, the machinery for such approximations is developed in §§8 under the rubric filtration; and 9 derives a metric ($d$-metric) on a base space of a filtration with respect to which the unit ball is compact.
The following two sections demonstrate the utility of that machinery by extending all the basic results to the separable setting. §10 shows that $\mathsf{CP}(\mathcal{H})$ is $d$-closed, that Kraus decompositions exist, and that they are extremal decompositions. It is perhaps surprising that existence of Kraus decomposition can be shown this way since the standard approach[11, 12] makes it seem a fairly deep result of $C^{*}$-algebra representation theory.
Finally, §11 tackles the GKSL problem in the separable setting, avoiding the $C^{*}$-algebraic tools deployed by Lindblad[2]. All the basic results of the finite-dimensional theory are extended: acceptability of generators (norm) continuously varying in the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ which coincides with the $L$-operators, and the existence of Lindblad parametrizers. In addition, we show that $\mathsf{cp}^{+}(\mathcal{H})$ and $\mathsf{cp}(\mathcal{H})$ are $d$-closed (hence, e.g. weak-operator closed).
A table of notations and terminology can be found in Appendix A.
@@ 2.3 Jamiołkowski transformation
2.3.1 A fundamental isomorphism: $\mathfrak{B}(()\mathcal{H},\mathcal{K})\cong\mathcal{K}\otimes\mathcal{H}^{*}$
@@
2.3.2 Jamiołkowski transformation defined
@@ 3.2.1 The trouble with $\mathsf{Mon}$
3.2.2 CP makes the world safe for tensor products
@@ 5 Markovian CP evolution and tangent geometry of $\mathsf{CP}(\mathcal{H})$
Here is a capsule summary, using some terminology that will be explained in the body.
  1. 1.
    In order for $\Lambda(t)$ to generate a right-differentiable Markovian CP evolution (generalizing CP semigroups),
    • it is necessary that $\Lambda(t)$ be in the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathfrak{B}(\mathcal{H})}$,
    • it is sufficient that $\Lambda(t)$ also be piecewise continuous.
  2. 2.
    The tangent space $\mathsf{cp}(\mathcal{H})$ is that part of the tangent cone consisting of operators tangent to $\mathsf{CP}(\mathcal{H})$ both “forward and backward”. These maps correspond to reversible evolutions.
  3. 3.
    $\mathsf{cp}^{+}(\mathcal{H})$ consists of the operators in $\mathsf{HP}(\mathcal{H})$ which can be parametrized (nonuniquely) by $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$ as $L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;]$.
  4. 4.
    Linear operators (Lindblad parametrizers) exist which (i) extract $L$-parametrizations, (ii) map $\mathsf{cp}^{+}(\mathcal{H})$ into $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$, and (iii) map the tangent cone $\mathsf{cp}^{+}(\mathcal{H})$ into $0\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. These provide a parametrization varying continuously with the generator.
  5. 5.
    Conventional parametrization results from Kraus decomposition of $\Psi$ and imposition of trace-preservation.
@@ 5.7.2 The concept
It is good to recall again at this point that, necessarily, if the first component of $\Delta\Lambda$ is CP, then $\Lambda$ is $\mathsf{cp}^{+}$.
@@ 7.4 Meanings of old notations in the separable setting
$\mathsf{Lind\,{\mathcal{H}}}=\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}},$ (7.9)
@@
8 Filtrations: sequences of finite-dimensional approximations
@@ 8.3 Generic notation
9 A filtration induces a complete, boundedly-precompact metric on its base space
@@
The $d$-topology is very weak. For instance, for $T\in\mathfrak{B}(\mathcal{H},\mathcal{K})$, $\mathscr{P}_{\alpha}T\mathscr{P}_{\alpha}$ amounts to a finite number of coordinates (linear functionals) of $Tv$ for a finite number of $v$ (basis of $\mathcal{H}_{\alpha}$). Thus, it is weaker than WOT. To put it another way, there are much fewer $d$-closed (or open) sets that WOT-closed, in some vague sense. Thus, it is gratifying to have the following criterion, which will be used in §10.2.
@@ 10 CP maps in separable setting
Item 1 means that if a sequence of CP maps converges in $d$-topology (or WOT, SOT, or norm), then the limit is also CP. The equivalence of item 2 above has its own interest, but is also crucial to item 3. The standard proof that Kraus decompositions exists rests on Stinespring’s theorem and a knowledge of the representations of the $C^{*}$-algebra $\mathfrak{B}(\mathcal{H})$. An explanation of how to use these results to show existence of Kraus decompositions can be found in Theorem 3.3 of [11] or Chapter 3 of [12]. Expositions of Stinespring’s theorem itself can be found in 34.7 of Ref. [23], Thm. IX.4.2 of Ref. [24], Thm. IV.3.6 of Ref. [25], or §9.2 of Ref. [26].
In comparison to the traditional method, which feels like a deus ex machina, the approach taken here is very pedestrian. We show that for any nonzero CP map $\Lambda$, we can subtract a $\bm{\theta}$}-map and have a CP map left over. Then we show that if one just keeps doing that (with some good choices), everything will converge as desired. It is worth pointing out that, even at the level of the result, existence of Kraus decompositions in the form of a sum is not obviously to be expected based on the finite-dimensional situation. From a convex analysis perspective, or considering the phenomenon of continuous spectrum, some sort of integral would not be surprising.
@@
10.1 Partial orders on $\mathsf{HP}(\mathcal{H})$
@@
10.2 $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is $d$-closed
@@
All this means is that if $\Lambda_{n}\precsim\Gamma_{n}$, $\Lambda_{n}\xrightarrow{d}\Lambda$, and $\Gamma_{n}\xrightarrow{d}\Gamma$, then $\Lambda\precsim\Gamma$. It follows immediately from Prop. 10.5 because what those relations mean is $\Gamma_{n}-\Lambda_{n}\in\mathsf{CP}(\mathcal{H},\mathcal{K})$. Conversely, Prop. 10.5 is also an immediate corollary of Prop. 10.6.
@@ 10.4 Extreme CP map = $\bm{\theta}$-map
Let $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ where $\mathcal{H}$, $\mathcal{K}$ are finite-dimensional, and let $\Lambda{\bm{\pi}}(h)$ have orthogonal decomposition $\Lambda{\bm{\pi}}(h)={\bm{\pi}}(k)+\cdots$. Then there exists a choice of $A_{i}$ in the Kraus decomposition $\Lambda=\sum_{i}{\bm{\theta}}(A_{i})$ such that
@@ 10.5 Order closedness of $\mathsf{CP}(\mathcal{H},\mathcal{K})$
Let $D$ be a $\leq$-directed set in $\mathsf{HP}(\mathcal{H},\mathcal{K})$, bounded above. Then $\sup D$ exists and $D$ converges to it in SOT. An analogous statement holds for a $\geq$-directed set.
@@
If $D$ is a bounded-above $\leq$-directed set contained in $\mathsf{CP}(\mathcal{H},\mathcal{K})$, then so is its supremum. An analogous statement holds for a $\geq$-directed set.
@@ 11.2 Projector juggling
$L\colon\mathsf{Lind\,{\mathcal{H}}}\rightarrow\mathsf{HP}(\mathcal{H})$ is defined by $L(\Psi,G,H)=\Psi-[G,\;\cdot\;]_{+}-[iH,\;\cdot\;]$, just as in finite dimensions. Since the formula is the same, $(L\circ\Delta^{\scriptstyle{min}}_{\alpha})\Gamma=\Gamma$ if $\Gamma\in\mathsf{HP}(\mathcal{H}_{\alpha})$. So, if we replace $\Gamma$ in the preceding equation by $\mathscr{P}_{\alpha}\Lambda$, we deduce
@@
11.3 Lindblad parametrizations and parametrizers exist
The main idea now is to show existence of Lindblad parametrizations and parametrizers by using the sequence $(\Delta^{\scriptstyle{min}}_{\alpha})_{\alpha}$ as some sort of approximation to a Lindblad parametrizer. For it to fulfill that role, we need the properties established in the first couple of subsections.
11.3.1 The first component of $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP when $\Lambda$ is $\mathsf{cp}^{+}$
Lemma 11.2.
If $\Lambda$ is $\mathsf{cp}^{+}$, then the first component of $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP.
Proof.
Because $\mathscr{P}_{\alpha}$ is a contraction, if $\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}}$, then $\mathscr{P}_{\alpha}\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H}_{\alpha})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}_{\alpha}}$. Therefore, as an element of $HP(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})$, $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ has CP first component. But it is still CP regarded as an element of $\mathsf{HP}(\mathcal{H})$. ∎
@@ 11.3.2 The sequence $(\Delta^{\scriptstyle{min}}_{\alpha})$ is an approximate right-inverse of $L$
$L\circ\Delta^{\scriptstyle{min}}_{\alpha}(\Lambda)\xrightarrow{SOT}\Lambda$ as $\alpha\to\infty$.
@@
We need the following calculation:
@@
Now, since (Lemma 11.1) $L$ and $\Delta^{\scriptstyle{min}}_{\alpha}$ are bounded uniformly in $\alpha$, it suffices to show that $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)\cdot A\rightarrow\Lambda A$ for $A$ in the set $\cup_{m}\mathcal{B}^{1}(\mathcal{H}_{m})$ dense in $\mathcal{B}^{1}(\mathcal{H})$. Thus, take AA in $\mathcal{B}^{1}(\mathcal{H}_{\beta})$. As soon as $\beta\leq\alpha$, $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)A=(\mathscr{P}_{\alpha}\Lambda)A$, by (11.10). But, $\mathscr{P}_{\alpha}\Lambda\xrightarrow{SOT}\Lambda$, and the conclusion follows immediately. ∎
@@ 11.3.3 Bringing in the $d$-metrics
Eventually, we must work with the $d$-metrics because that is the only topology we have with respect to which every bounded sequence has cluster points.